OpenStudy (anonymous):

Log 2e^-2t + 5e^-t = 3, determine the value of t. Need someone to double check. I got t=3, is it correct? Thanks !!

OpenStudy (cggurumanjunath):

can u type the equation ?

OpenStudy (ankit042):

One way to check the solution is by putting back the value of t in equation. If solution is correct you will get zero after putting value of t. actually e^-t =-3 or e^-t =1/2 now you can solve to get t

OpenStudy (anonymous):

\[2e ^{-2t}+5e ^{-t}-3=0\] \[Put e ^{-t}=x,squaring,e ^{-2t}=x ^{2}\] \[2x ^{2}+5x+3=0\] make factors and find the value of x i.e., \[e ^{-t}\]

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

yea my equation is 2x^2+5x-3=0 not "+3" am I right? the answer I got is t= 1/3 right ?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[ 2x^2+5x-3=0\] \[(x+3)(2x-1)=0\] so \[x=-3\] or \[x=\frac{1}{2}\] but you are still not done here

OpenStudy (anonymous):

because your job was not to solve for \(x\) but for \(t\) if \(x=e^{-t}\) this is where the log comes in

OpenStudy (anonymous):

\[e^{-t}=-3\] has no solution as \(e\) to any power is positive

OpenStudy (anonymous):

\[e^{-t}=\frac{1}{2}\] can you solve that?

OpenStudy (anonymous):

1/e^t = 1/2 e^t = 2 t ln e= ln2 t = 2?

OpenStudy (anonymous):

last line is wrong

OpenStudy (anonymous):

\[e^{-t}=\frac{1}{2}\iff -t=\ln(\frac{1}{2})=-\ln(2)\] so \[t=\ln(2)\]

OpenStudy (anonymous):

or if you prefer \[e^t=2\iff t=\ln(2)\] same thinig

OpenStudy (anonymous):

oh so it 0.69!

OpenStudy (anonymous):

as a decimal approximation, yes

OpenStudy (anonymous):

thank you!!